OpenStudy (anonymous):

Pls help: jackie usually mows his lawn in 3 hours marilyn can mow the same yard in 5 hours how much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{1}{5}= \frac{1}{x} \]x=15/8 hours or 1:52.5

OpenStudy (rosey):

15/8 u should b a fan of me :)